Not sure if I have this in the right category. Thanks for your help.
P(4,0), Q(6,-8)
use point-slope form to write the equation of the line passing through the two given points. Write each equation in slope-intercept form. Thanks!

Answer by checkley71(8403)

Y=mX+b
m=(-8-0)/6-4)
m=-8/2
m=-4
Y=-4X+b
SUBSTITUTING ONE OF THE POINTS IN THIS EQUATION WE FIND THE INTERCEPT (b)
0=4*-4+b
0=-16+b
b=16
THUS THE EQUATION IS
Y=-4X-16